# The Unreasonable Ineffectiveness of the Deeper Layers

### v1

@@ A.2 Evaluation details
  • We follow the formatting suggested in the original reference Hendrycks et al. (2020) without further prompt engineering.
@@
  • For our experiments, we use 0 few-shot examples; our results and analysis are robust to this choice, cf. Figure 7.
@@
  • For our experiments, we use 0 few-shot examples.
@@
  • In Figure 3 we normalized the loss to facilitate fair comparison across model families that employ different vocab sizes: to normalize, we divided by $\log V$, where $V$ is the per-model vocab size (listed in a table in §A.1). This, $\log V$, corresponds to the loss of sampling tokens uniformly, which naturally sets the scale for a given model.
Appendix B Ablations
Here we detail ablations of various hyperparameters: prompting (§B.1), finetuning seed (§B.2), LoRA rank (§B.3). Qualitatively, the results of the paper are quite robust to the variation of any of these.
B.1 Prompting
@@ B.3 LoRA rank
  • Begin with rank 64, following the QLoRA setup (see, e.g. Appendix B.2 of Ref. (Dettmers et al., 2023)).
@@
In practice, this led to us using rank 64 for every model with the exceptions of Mistral-7B, with rank 4, Llama-2-7B, with rank 2, and Llama-2-70B, with rank 8. (To review this same information in tabular form, see the second Table in §A.1.) Figure 9 displays the sweeps over MMLU accuracy supporting these choices for Mistral-7B (bottom left panel), Llama-2-7B (bottom middle panel), and Llama-2-70B (top right panel): overall, while the LoRA rank does not have a significant impact on the qualitative behavior of the healed model, decreasing the LoRA rank generally improves performance. In the top left and middle panels of Figure 9, we show corresponding sweeps for Mistral-7B (top) and Llama-2-7B (middle) using the similarity-informed pruning strategy: we see that for this pruning method both models are much more robust, though rank 2 is still the top performing rank for Llama-2-7B.
@@
The characteristic improvement of MMLU accuracy with decreasing LoRA rank – even for extremely low ranks(!) – deserves an explanation. One possibility is that lowering the LoRA rank can better regularize finetuning against overfitting. In particular, astute readers may have been surprised at the discussion of peak learning rates in §A.1: models were finetuned with the same peak used in pretraining; a “large” LoRA rank of 64 introduces a number of additional parameters that may overfit to C4. This overfitting would certainly be harmful, since the actual pretraining datasets for the models we consider are (a) unknown to us, and (b), likely to be of significantly higher quality than C4.

### v2

@@ B.2 Evaluation details
  • We follow the formatting suggested in the original reference (Hendrycks et al., 2020) without further prompt engineering.
@@
  • For our experiments, we use 00 few-shot examples; our results and analysis are robust to this choice, cf. Figure 8.
@@
  • For our experiments, we use 00 few-shot examples.
@@
  • In Figure 3 we normalized the loss to facilitate fair comparison across model families that employ different vocab sizes: to normalize, we divided by $\log V$, where $V$ is the per-model vocab size (listed in a table in §B.1). This, $\log V$, corresponds to the loss of sampling tokens uniformly, which naturally sets the scale for a given model.
Appendix C Ablations
Here we detail various ablations: prompting (§C.1), finetuning seed (§C.2), LoRA rank (§C.3), other pruning strategies (§C.4). Qualitatively, the results of the paper are quite robust to the variation of any of these.
C.1 Prompting
@@ C.3 LoRA rank
  • Begin with rank 64, following the QLoRA setup (see, e.g. Appendix B.2 of Dettmers et al. (2023)).
@@
In practice, this led to us using rank 64 for every model with the exceptions of Mistral-7B, with rank 4, Llama-2-7B, with rank 2, and Llama-2-70B, with rank 8. (To review this same information in tabular form, see the second Table in §B.1.) Figure 10 displays the sweeps over MMLU accuracy supporting these choices for Mistral-7B (bottom left panel), Llama-2-7B (bottom middle panel), and Llama-2-70B (top right panel): overall, while the LoRA rank does not have a significant impact on the qualitative behavior of the healed model, decreasing the LoRA rank generally improves performance. In the top left and middle panels of Figure 10, we show corresponding sweeps for Mistral-7B (top) and Llama-2-7B (middle) using the similarity-informed pruning strategy: we see that for this pruning method both models are much more robust, though rank 2 is still the top performing rank for Llama-2-7B.
@@
The characteristic improvement of MMLU accuracy with decreasing LoRA rank – even for extremely low ranks(!) – deserves an explanation. One possibility is that lowering the LoRA rank can better regularize finetuning against overfitting. In particular, astute readers may have been surprised at the discussion of peak learning rates in §B.1: models were finetuned with the same peak used in pretraining; a “large” LoRA rank of 64 introduces a number of additional parameters that may overfit to C4. This overfitting would certainly be harmful, since the actual pretraining datasets for the models we consider are (a) unknown to us, and (b), likely to be of significantly higher quality than C4.
@@
C.4 Other pruning strategies
Here we study how the similarity-informed pruning strategy (§ 3.2) compares to other layer-pruning baselines: specifically, we contrast with pruning random layers and pruning shallow layers. In Figure 11, we observe that the similarity-informed strategy from the main text outperforms both of these other strategies on an MMLU evaluation of Llama-7B.
Figure 11: Comparison of the similarity-informed pruning strategy (blue) to random-layer pruning (orange) and shallow-layer pruning (green) on MMLU accuracy, with Llama-2 7B and LoRA rank 64. The similarity-informed pruning strategy clearly outperforms these baselines.